You wish to create a complete portfolio with 12 % in T-Bills and the rest in a risky portfolio. The risky portfolio is has 69 % in stocks and the rest in bonds.

E(r) Std Beta
T-Bills 1.2 %
Stocks 22.8 % 28% 1,4
Bonds 6.4 % 6% 0,2

What is the complete portfolio expected return? Please convert your answer to percentage points and use two decimal points in your answer. If your answer for example is 0.0592 = 5.92 %, then please write 5.92 as your answer.

Also remember that you need to use a dot and not a comma to represent your decimal separator.
Answer #1

Expected return of a portfolio is weighted average of return of individual portfolio components.

E(R) = W1 * r1 + W2 * r2 + W3 * r3

W1 (weight in Treasury) = 12%

Weight in risky portfolio = 1 - 12% = 88%

W2 (weight in stocks) = 88% * 69% = 60.72%

W3 (weight in bonds) = 88% * (1 - 69%) = 27.28%

E(R) = 12% * 1.2% + 60.72% * 22.8% + 27.28% * 6.4%

E(R) = 0.14% + 13.84% + 1.75%

E(R) = 15.73%

Hence, answer should be entered as 15.73